OpenStudy (anonymous):

The perimeter of a rectangle is 72 centimeters. The base is 2 times the height. What are the dimensions of the rectangle? a. h = 9 cm, b = 18 cm c. h = 12 cm, b = 24 cm b. h = 19 cm, b = 38 cm d. h = 10 cm, b = 20 cm

OpenStudy (anonymous):

2(l+b) = perimeter 2(x+2x)=72

OpenStudy (anonymous):

Translate the expression into a word phrase. k + 12 A. a number less than 12 B. 12 more than a number C. the product of a number and 12 D. the quotient of a number and 12

OpenStudy (anonymous):

2L+2W=72 cm L=2W 2(2W)+2W=72 4W+2W=72 6W=72 W=12 cm L=2W L=2(12) L=24 cm So it's letter C.